OpenStudy (anonymous):

lim x approaches 0 for f(x)=(1-x)/(x(1+xk)^(n-1))

OpenStudy (anonymous):

k is positive

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{1-x}{x(1+xk)^{n-1}}\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

give me a second to think

OpenStudy (anonymous):

wolfram is giving me +/- infinity

myininaya (myininaya):

what about n

OpenStudy (anonymous):

I'll type out the rest of the word problem

OpenStudy (anonymous):

In an article on the local clustering of cell surface receptors, the researcher analyzed the limits of the function: (The function I put) X is the concentration of free receptors, n is the number of functional groups of cells, and k is a positive constant The function is also equal to C which is the concentration of free ligand in the medium

OpenStudy (anonymous):

i think its supposed to represent a previous amount

myininaya (myininaya):

yeah i don't know about all that stuff i'm sorry astro

myininaya (myininaya):

what is this called?

myininaya (myininaya):

i can try to find someone who knows about whatever this is

OpenStudy (anonymous):

no problem- its just studying limits for now based around cell surface receptors

myininaya (myininaya):

and we are assuming maybe n>1?

OpenStudy (anonymous):

That would make sense

myininaya (myininaya):

ok i think i got it one sec

myininaya (myininaya):

\[\lim_{x \rightarrow 0^+}\frac{1-x}{x(1+xk)^{n-1}}=\infty \] since 1-x is negative for any values approaching 0 from the right but both the factors of the bottom are positive since k>0 we know nothing can be done to force f to be continuous at x=0 so we know from the right it approaches +infinity and from the left it approaches negative infinity since 1-x is positive at x approaches 0 from the left

myininaya (myininaya):

as approaches*

OpenStudy (anonymous):

Thank you, that makes sense to me

myininaya (myininaya):

cool! :)

myininaya (myininaya):

i'm glad i could help you with the math part lol